# Magnetic Anti-de Sitter Wormholes as seeds for Higgs Inflation

### v1

@@ IV Discussion
In this Letter, we proposed that certain magnetic Euclidean (half) wormhole saddles can catalyze the onset of inflation, while at the same time being embeddable in a model of Higgs inflation containing only experimentally observed SM fields. This constitutes an important development compared to Betzios:2024oli , since the models we consider in the present work lead to predictions that are in line with current observational data.
The most interesting future direction is to analyze in great detail the quantum effective potential for the Higgs, assuming various scenarios for particle physics at the inflationary scale and above (for example the presence or not of Supersymmetry and Grand-Unification). Our wormhole backgrounds exist only if the electroweak vacuum is metastable, the energy window where the effective potential turns negative being higher than the inflationary scale. This allows us to place stringent bounds on particle masses and perhaps other SM parameters, by solving the RG equations for each such scenario and therefore to solidify or rule out our proposal. The presence of primordial magnetic fields in our model also offers a potential window into understanding the origin of the large scale magnetization of our Universe, albeit it suffers from the same inflationary dilution issues that plague similar proposals in the literature.
Finally, due to their Anti-de Sitter asymptotics, such Euclidean wormhole saddles Maldacena:2004rf ; Betzios:2019rds ; Betzios:2021fnm ; VanRaamsdonk:2021qgv ; Antonini:2022ptt ; Betzios:2024oli could descend from a fundamental UV complete construction (String Theory or the AdS/CFT correspondence). At a more conservative level, these Euclidean saddles are an effective proxy for the expected complicated chaotic dynamics of the early universe, allowing us to perform consistent semi-classical computations, in analogy to studies of quantum mechanical black holes Almheiri:2020cfm .
V acknowledgments
@@
  • According to ParticleDataGroup:2024cfk the average value of the Higgs mass is $m_{H}=125.20\pm 0.11\,{\rm GeV}$, while the top quark mass computed from direct measurements is $m_{t}=172.57\pm 0.29\,{\rm GeV}$. The stability is very sensitive to $m_{t}$. The Higgs self coupling $\lambda(\mu)$ following from SM RG equations decreases up to some scale and then increases. Whether it remains positive or not depends on the Higgs mass and the top Yukawa coupling $y_{t}$ at each RG scale $\mu$. Within the present uncertainties there is a case of both stability, metastability or instability Masina:2012tz ; Espinosa:2015qea ; Bednyakov:2015sca . Based on the central values provided by ParticleDataGroup:2024cfk we are in the metastable region. Gravitational effects on vacuum stability were initially studied in Coleman:1980aw , while the impact of gravity on the stability of the SM Higgs has been discussed in Isidori:2007vm ; Salvio:2016mvj ; Branchina:2016bws ; Markkanen:2018pdo ; Rajantie:2016hkj ; Espinosa:2020qtq , as well as in alternative formulations of gravity in Gialamas:2022gxv ; Gialamas:2023emn .
  • We also have the following cutoff scales for the effective field theory of Higgs inflation: The higher cutoff is the Planck scale $\Lambda_{\rm UV}\leq M_{\rm Pl}$. The value of the cutoff in general depends on the VEV/value of the Higgs boson $\phi$. Above the EW vacuum the cutoff is Burgess:2009ea ; Barbon:2009ya $\Lambda_{\rm ew}=M_{\rm Pl}/\xi$. In the inflationary regime the cutoff is $\Lambda_{\rm inf}=M_{\rm Pl}/\sqrt{\xi}$, which for values of $\xi<10^{3}$ that we are interested in, is much higher than the energies of the processes of interest. In the case of metastable Higgs inflation the value of $\xi$ can be even lower than that Bezrukov:2014ipa (see also McDonald:2023qzz ), raising the value of the cutoff even further, giving further credence to our scenario.
@@
The scalar field part of the action (II) of the main text can be expressed in the Einstein frame through a Weyl rescaling of the form
@@
Integrating the previous equation we obtain Garcia-Bellido:2008ycs
@@
In this model at tree level, the inflationary observables (for $N_{\star}\gg 1$) come out to be Bezrukov:2007ep
@@
Observations of the cosmic microwave background (CMB) place significant constraints on inflationary predictions, as shown in Planck:2018jri ; BICEP:2021xfz . The latest combined data from Planck, BICEP/Keck, and BAO have set the following limits on observable values at the pivot scale $k_{\star}=0.05\,{\rm Mpc}^{-1}$:
@@
An analysis of Higgs non-minimal $\xi$-inflation must also include the renormalization group (RG) running of the couplings and loop corrections to the (effective) action. Under RG, the minimal set of higher-dimensional operators to be included on top of the tree-level action is generated by the theory itself via radiative corrections. Since our model of Einstein gravity coupled to the SM constitutes a non-renormalizable theory, the cancellation of the loop divergences stemming from the original action requires the inclusion of an infinite set of counterterms. We can therefore threat this model only as an effective field theory (EFT). The outcome of the subtraction procedure of divergences depends on the renormalization scheme, with different choices corresponding to different assumptions about the ultraviolet (UV) completion of the model. In any RG-scheme, the best prescription is to chose a scale $\mu$ such that the effect of the logarithms is minimised as possible (so in general the scale will also depend on the background field values i.e. $\mu(\phi)$). Moreover this choice of scheme is not frame independent, but transforms under Weyl transformations (2) as $\mu\rightarrow\mu/\Omega$.
In our context of Higgs inflation, there exist two basic inequivalent renormalization prescriptions/schemes depending on whether quantum corrections to the potential and effective action are computed setting a fixed, field independent cutoff scale in the Einstein frame (called prescription I) Bezrukov:2007ep ; Bezrukov:2008ej ; Bezrukov:2014ipa , or in the Jordan frame (called prescription II) Barvinsky:2008ia ; Shaposhnikov:2008xi ; Barvinsky:2008ia ; DeSimone:2008ei ; Barvinsky:2009fy ; Barvinsky:2009ii . These correspond to different UV-completions of the model. The first prescription is connected to the idea of obtaining quantum scale invariance at high energies, while the second is motivated by using the standard RG prescription for computing loop corrections and measuring physical distances and scales with respect to a fixed Planck cutoff.
We should also add at this point that there exists a third prescription (III) Markkanen:2018bfx in which the RG-scale $\mu$ depends also on the curvature $R$, which can potentially become important if the size of the early universe is sufficiently small and approaches the Planck scale. In the next section, we prove that in our model it is safe to neglect any such curvature dependence to define the renormalization point/substraction scheme, since the relevant wormhole backgrounds that we consider have a curvature scale that is found to be at least nine orders smaller than $M_{\rm Pl}^{2}$.
@@
and initial conditions at $\mu_{0}$. The renormalization point is chosen such that one minimizes the logarithms as much as possible. At this point, and in line to our previous discussion, one should keep in mind that the precise shape of the effective potential is gauge and RG scheme dependent. Nevertheless both the extrema and the value of the potential at them are gauge independent as demanded by the Nielsen identity Nielsen:1975fs (this provides us with crucial information about various physical couplings and observables at the inflationary plateau and electroweak scale).
@@
in terms of Yukawa couplings ($y_{i}$) for each fermion of the SM. The SM beta functions at one-loop are then given by Buttazzo:2013uya
@@
When including gravity at one-loop order, one needs to include the following additional terms in the purely gravitational part of the Euclidean effective action tHooft:1974toh
@@
Collecting the contributions of the various SM fields, the gravitational part of the $\beta$-functions is (this set together with eqns. (RG Improved effective potential) is exhaustive to one-loop order Markkanen:2018bfx , see also Deser:1974cy ; Deser:1974cz ; Deser:1974xq , for results using dimensional regularization)
@@
Assuming that quantum fluctuations of the fields do not change the symmetry of the background metric, in our FLRW type of background (7), the quadratic in curvature terms combine to give
@@
We shall now describe the properties of the one-loop RGE improved effective potential $V_{\rm RGE}(\phi;g_{i};\mu)$ for the scalar (Higgs) field. In order to do so it is convenient to introduce the Heat kernel coefficients Birrell:1982ix ; Parker:2009uva ; Markkanen:2018bfx
@@
in terms of coefficients that depend on the running couplings of the theory (their explicit form can be found in Markkanen:2018bfx ).
@@
Table 1: The 1-loop contributions with tree-level couplings to the Higgs. $\Psi$ stands for $W^{\pm}$ and $Z^{0}$ bosons, the 6 quarks q, the 3 charged leptons $l$, the Higgs $h$. The Goldstone bosons are $\chi_{W}$ and $\chi_{Z}$ and ghosts $c_{W}$ and $c_{Z}$. The masses may be found Markkanen:2018bfx . (This table can also be found in Markkanen:2018bfx ).
Our analysis proves that in our model, up to a tremendous degree of accuracy ($\sim 10^{-8}$), that holds up to the inflationary scale and above, we can compute the loop corrected effective potential for the scalar Higgs using flat space RG results. This has already been performed in various works Bezrukov:2009db ; Barvinsky:2009fy , using the type I and II schemes Bezrukov:2007ep ; Bezrukov:2008ej ; Bezrukov:2014ipa ; Barvinsky:2008ia ; Shaposhnikov:2008xi ; Barvinsky:2008ia ; DeSimone:2008ei ; Barvinsky:2009ii . Our work shows that the best renormalization scheme from an EFT perspective is the one that reduces the logarithms as much as possible, so that in general $\mu^{2}=A\phi_{cl}^{2}+CM_{\rm Pl}^{2}$. In contrast with both schemes I, II in the literature, the precise numerical value of $A,C$ is properly fixed by minimizing logarithms at the scales of interest (i.e. inflationary scale $\mu_{\star}$), without the need to assume certain features for the UV completion of our EFT model. This result is reasonable since EFTs are constructed such that they are as predictive as possible at the chosen physical scales of interest. An analysis of the RGE improved potential for a range of parameters of the SM is presented in Masina:2024ybn . The detailed investigation of the parameter space of the SM (and its possible supersymmetric or GUT extensions at the inflationary or higher energy scales) to ensure alignment with our model will be the subject of future work.

### v2

@@ IV Discussion
In this paper, we proposed that certain magnetic Euclidean (half) wormhole saddles can catalyze the onset of inflation, while at the same time being embeddable in a model of Higgs inflation containing only experimentally observed SM fields. This constitutes an important development compared to Betzios and Papadoulaki (2024), since the models we consider in the present work lead to predictions that are in line with current observational data.
The most interesting future direction is to analyze in great detail the quantum effective potential for the Higgs, assuming various scenarios for particle physics at the inflationary scale and above (for example the presence or not of Supersymmetry and Grand-Unification). Our wormhole backgrounds exist only if the electroweak vacuum is metastable, the energy window where the effective potential turns negative being higher than the inflationary scale. RG corrections, could also potentially leave intermediate scale imprints affecting the detailed numbers of observational quantities, such as the spectral index. This allows us to place stringent bounds on particle masses and perhaps other SM parameters, by solving the RG equations for each such scenario and therefore to solidify or rule out our proposal. The presence of primordial magnetic fields in our model also offers a potential window into understanding the origin of the large scale magnetization of our Universe, albeit it suffers from the same inflationary dilution issues that plague similar proposals in the literature.
Finally, due to their Anti-de Sitter asymptotics, such Euclidean wormhole saddles Maldacena and Maoz (2004); Betzios et al. (2019, 2022); Van Raamsdonk (2022); Antonini et al. (2023); Betzios and Papadoulaki (2024) could descend from a fundamental UV complete construction (String Theory or the AdS/CFT correspondence). At a more conservative level, these Euclidean saddles are an effective proxy for the expected complicated chaotic dynamics of the early universe, allowing us to perform consistent semi-classical computations, in analogy to studies of quantum mechanical black holes Almheiri et al. (2021).
Acknowledgements.
We wish to acknowledge useful discussions with several colleagues in conferences and seminars where we have presented our work. We also wish to thank the anonymous referees for their comments that helped us improve the quality and clarity of the manuscript.
@@ Acknowledgements.
  • According to Navas et al. (2024) the average value of the Higgs mass is $m_{H}=125.20\pm 0.11\,{\rm GeV}$, while the top quark mass computed from direct measurements is $m_{t}=172.57\pm 0.29\,{\rm GeV}$. The stability is very sensitive to $m_{t}$. The Higgs self coupling $\lambda(\mu)$ following from SM RG equations decreases up to some scale and then increases. Whether it remains positive or not depends on the Higgs mass and the top Yukawa coupling $y_{t}$ at each RG scale $\mu$. Within the present uncertainties there is a case of both stability, metastability or instability Masina (2013); Espinosa et al. (2015); Bednyakov et al. (2015). Based on the central values provided by Navas et al. (2024) we are in the metastable region. Gravitational effects on vacuum stability were initially studied in Coleman and De Luccia (1980), while the impact of gravity on the stability of the SM Higgs has been discussed in Isidori et al. (2008); Salvio et al. (2016); Branchina et al. (2016); Markkanen et al. (2018a); Rajantie and Stopyra (2017); Espinosa (2020), as well as in alternative formulations of gravity in Gialamas et al. (2023); Gialamas and Veermäe (2023).
  • We also have the following cutoff scales for the effective field theory of Higgs inflation: The higher cutoff is the Planck scale $\Lambda_{\rm UV}\leq M_{\rm Pl}$. The value of the cutoff in general depends on the VEV/value of the Higgs boson $\phi$. Above the EW vacuum the cutoff is Burgess et al. (2009); Barbon and Espinosa (2009) $\Lambda_{\rm ew}=M_{\rm Pl}/\xi$. In the inflationary regime the cutoff is $\Lambda_{\rm inf}=M_{\rm Pl}/\sqrt{\xi}$, which for values of $\xi<10^{3}$ that we are interested in, is much higher than the energies of the processes of interest. In the case of metastable Higgs inflation the value of $\xi$ can be even lower than that Bezrukov et al. (2015) (see also McDonald (2024)), raising the value of the cutoff even further, giving further credence to our scenario.
@@
The scalar field part of the action (1) of the main text can be expressed in the Einstein frame through a Weyl rescaling of the form
@@
Integrating the previous equation we obtain Garcia-Bellido et al. (2009)
@@
In this model at tree level, the inflationary observables (for $N_{\star}\gg 1$) come out to be Bezrukov and Shaposhnikov (2008)
@@
Observations of the cosmic microwave background (CMB) place significant constraints on inflationary predictions, as shown in Akrami et al. (2020); Ade et al. (2021). The latest combined data from Planck, BICEP/Keck, and BAO have set the following limits on observable values at the pivot scale $k_{\star}=0.05\,{\rm Mpc}^{-1}$:
@@
An analysis of Higgs non-minimal $\xi$-inflation must also include the renormalization group (RG) running of the couplings and loop corrections to the (effective) action. Under RG, the minimal set of higher-dimensional operators to be included on top of the tree-level action is generated by the theory itself via radiative corrections. Since our model of Einstein gravity coupled to the SM constitutes a non-renormalizable theory, the cancellation of the loop divergences stemming from the original action requires the inclusion of an infinite set of counterterms. We can therefore threat this model only as an effective field theory (EFT). The outcome of the subtraction procedure of divergences depends on the renormalization scheme, with different choices corresponding to different assumptions about the ultraviolet (UV) completion of the model. In any RG-scheme, the best prescription is to chose a scale $\mu$ such that the effect of the logarithms is minimised as possible (so in general the scale will also depend on the background field values i.e. $\mu(\phi)$). Moreover this choice of scheme is not frame independent, but transforms under Weyl transformations (16) as $\mu\rightarrow\mu/\Omega$. In our context of Higgs inflation, there exist two basic inequivalent renormalization prescriptions/schemes depending on whether quantum corrections to the potential and effective action are computed setting a fixed, field independent cutoff scale in the Einstein frame (called prescription I) Bezrukov and Shaposhnikov (2008); Bezrukov et al. (2009, 2015), or in the Jordan frame (called prescription II) Barvinsky et al. (2008); Shaposhnikov and Zenhausern (2009); Barvinsky et al. (2008); De Simone et al. (2009); Barvinsky et al. (2009, 2012). These correspond to different UV-completions of the model. The first prescription is connected to the idea of obtaining quantum scale invariance at high energies, while the second is motivated by using the standard RG prescription for computing loop corrections and measuring physical distances and scales with respect to a fixed Planck cutoff.
We should also add at this point that there exists a third prescription (III) Markkanen et al. (2018b) in which the RG-scale $\mu$ depends also on the curvature $R$, which can potentially become important if the size of the early universe is sufficiently small and approaches the Planck scale. In the next section, we prove that in our model it is safe to neglect any such curvature dependence to define the renormalization point/substraction scheme, since the relevant wormhole backgrounds that we consider have a curvature scale that is found to be at least nine orders smaller than $M_{\rm Pl}^{2}$.
@@
and initial conditions at $\mu_{0}$. The renormalization point is chosen such that one minimizes the logarithms as much as possible. At this point, and in line to our previous discussion, one should keep in mind that the precise shape of the effective potential is gauge and RG scheme dependent. Nevertheless both the extrema and the value of the potential at them are gauge independent as demanded by the Nielsen identity Nielsen (1975) (this provides us with crucial information about various physical couplings and observables at the inflationary plateau and electroweak scale).
@@
in terms of Yukawa couplings ($y_{i}$) for each fermion of the SM. The SM beta functions at one-loop are then given by Buttazzo et al. (2013)
@@
When including gravity at one-loop order, one needs to include the following additional terms in the purely gravitational part of the Euclidean effective action ’t Hooft and Veltman (1974)
@@
Collecting the contributions of the various SM fields, the gravitational part of the $\beta$-functions is (this set together with eqns. (RG Improved effective potential) is exhaustive to one-loop order Markkanen et al. (2018b), see also Deser and van Nieuwenhuizen (1974a, b); Deser et al. (1974), for results using dimensional regularization)
@@
Assuming that quantum fluctuations of the fields do not change the symmetry of the background metric, in our FLRW type of background (eq. (7) in the main text), the quadratic in curvature terms combine to give
@@
We shall now describe the properties of the one-loop RGE improved effective potential $V_{\rm RGE}(\phi;g_{i};\mu)$ for the scalar (Higgs) field. In order to do so it is convenient to introduce the Heat kernel coefficients Birrell and Davies (1982); Parker and Toms (2009); Markkanen et al. (2018b)
@@
in terms of coefficients that depend on the running couplings of the theory (their explicit form can be found in Markkanen et al. (2018b)).
@@
Table 1: The 1-loop contributions with tree-level couplings to the Higgs. $\Psi$ stands for $W^{\pm}$ and $Z^{0}$ bosons, the 6 quarks q, the 3 charged leptons $l$, the Higgs $h$. The Goldstone bosons are $\chi_{W}$ and $\chi_{Z}$ and ghosts $c_{W}$ and $c_{Z}$. The masses may be found Markkanen et al. (2018b). (This table can also be found in Markkanen et al. (2018b)).
Our analysis proves that in our model, up to a tremendous degree of accuracy ($\sim 10^{-8}$), that holds up to the inflationary scale and above, we can compute the loop corrected effective potential for the scalar Higgs using flat space RG results. This has already been performed in various works Bezrukov and Shaposhnikov (2009); Barvinsky et al. (2009), using the type I and II schemes Bezrukov and Shaposhnikov (2008); Bezrukov et al. (2009, 2015); Barvinsky et al. (2008); Shaposhnikov and Zenhausern (2009); Barvinsky et al. (2008); De Simone et al. (2009); Barvinsky et al. (2012). Our work shows that the best renormalization scheme from an EFT perspective is the one that reduces the logarithms as much as possible, so that in general $\mu^{2}=A\phi_{cl}^{2}+CM_{\rm Pl}^{2}$. In contrast with both schemes I, II in the literature, the precise numerical value of $A,C$ is properly fixed by minimizing logarithms at the scales of interest (i.e. inflationary scale $\mu_{\star}$), without the need to assume certain features for the UV completion of our EFT model. This result is reasonable since EFTs are constructed such that they are as predictive as possible at the chosen physical scales of interest. An analysis of the RGE improved potential for a range of parameters of the SM is presented in Masina and Quiros (2024). The detailed investigation of the parameter space of the SM (and its possible supersymmetric or GUT extensions at the inflationary or higher energy scales) to ensure alignment with our model will be the subject of future work.